solve the equation


 


cos3X sin3(cube)X + sin3X cos3(cube)X  =  0

Grade:11

1 Answers

suryakanth AskiitiansExpert-IITB
13 years ago

Dear somesh,

we know that,

cos^3x=(cos3x+3cosx)/4 and  sin^3x=(3sinx-sin3x)/4                    [since cos3x=4cos^x-3cosx]

substuting in the given equation

cos3X sin3(cube)X + sin3X cos3(cube)X  =  0

cos3x{(3sinx-sin3x)/4}+sin3x{(cos3x+3cosx)/4} = 0

{3sinxcos3x-cos3xsin3x+sin3xcos3x+3cosxsin3x}/4 = 0

3/4(sinxcos3x+cosxsin3x) = 0

sin4x = 0

Hence x=n∏/4

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum..